# Re-visiting Skip-Gram Negative Sampling: Dimension Regularization for More Efficient Dissimilarity Preservation in Graph Embeddings

### v1

@@ 3.1. Instantiation for LINE and node2vec
LINE and node2vec differ in their implementation of the random-walk generation function. For LINE, the function simply returns the edge set $E$. For node2vec, the function returns the set of all node and neighbor pairs ($\left\{i,j|S_{ij}=1\right\}$) where neighbors of $i$ are nodes encountered on a biased random walk starting at $i$. The parameters $p,q$ control the bias of the random walk, as specified in Grover and Leskovec (2016).
4. Evaluation
In this section, we present our evaluation of augmented LINE and augmented node2vec where SGNS has been replaced with dimension regularization via Algorithm 1. We show that on a variety of real-world networks, the augmented algorithms achieve better link-prediction results while also reducing training runtime. We also compare the augmented algorithms to baselines in which the negative function has been removed completely; while these baselines perform well on real-world networks, we show that they rely on low levels of connectivity to sustain link-prediction performance.
4.1. Evaluation Methodology
We evaluate the augmented algorithms on a downstream link-prediction task where the two evaluation metrics are Precision@$k$ and Transductive Precision@$k$. Both are detailed below:
Precision. (In-Sample)
After training an algorithm on a graph $G$ and obtaining an embedding matrix $X$, the precision@$k$ for a single node $i$ is the fraction of the top-k nodes most similar to node $i$ that are connected to $i$. Similarity between two nodes is defined as the dot product of their embeddings. The precision for the whole graph is the average node precision for all nodes with degree at least $k$. We refer to precision as in-sample precision since it is a measure of how much the embeddings overfit to the edges they are trained on. In all of our experiments, we set $k=10$.
Transductive Precision. (Out-of-Sample)
We also measure the link-prediction precision for the embeddings on a hold-out set of edges. Before training the embeddings, we uniformly randomly split $E$ into $E_{\text{train}}$ and $E_{\text{test}}$ following a $70-30$ split. Let $V^{\prime}$ be the set of nodes in the largest connected component induced by $E_{\text{train}}$. Then, we define train and test graphs $G_{\text{train}}=\left(V^{\prime},E_{\text{train}}\right)$ and $G_{\text{test}}=\left(V^{\prime},E_{\text{test}}\right)$. The transductive precision is the precision of embeddings trained on $G_{\text{train}}$ but evaluated on $G_{\text{test}}$. Similar to the previous section, we refer to transductive precision as out-of-sample precision because it is a measure of how much embeddings generalize to unseen edges.
Baselines.
We compare the augmented algorithms against (unaltered) LINE and node2vec, which we call the Vanilla baselines. In addition, we compare against baselines in which the negative function has been removed entirely, which we call the Positive Only baselines. We use these baselines to demonstrate that the embeddings collapse in practice without a negative function. We also analyze both the unnormalized and normalized Positive Only embeddings as past work has shown that normalizing embeddings improves downstream performance (Xu and Durrett, 2018).
Hyperparameters.
@@ 4.1. Evaluation Methodology
Datasets.
The datasets used for evaluation are listed in Table 2. We chose five real-world datasets spanning a variety of domains (Leskovec and Krevl, 2014). For all of the graphs, if there are multiple connected components we keep only the largest. The statistics in Table 2 characterize the largest connected component.
Table 2. Graph datasets used in our study.
Graph Type $n$ $m$
Wikipedia Language 4.8K 93K
Facebook Social 4.0K 88K
Protein-Protein Biological 3.9K 39K
ca-HepTh Citation 8.6K 25K
LastFM Social 7.6K 28K
4.2. Results
Our results answer three fundamental questions:
Q1: How well do the augmented algorithms perform on link prediction tasks relative to the original algorithms?
Figure 3 shows that the augmented embeddings, on average, achieve higher out-of-sample transductive precision values than their vanilla counterparts. In contrast, Figure 4 shows that augmented embeddings achieve lower in-sample precision values. The figures suggest that the augmented embeddings better generalize to unseen edges, whereas the vanilla embeddings overfit to the training dataset and struggle to generalize well. A notable case is the Wikipedia graph, where replacing SGNS with dimension regularization yields a $12$-fold increase in transductive precision. The improved performance of dimension regularization can be attributed to reduced overfitting on the training dataset, as witnessed by the fact that on average, in-sample Precision@$k$ decreased by $36.7$% for LINE and $15.4$% for node2vec across the five graphs.
The Positive Only baselines perform well both in-sample and out-of-sample at times, but do not do so consistently, as seen in the cases of LastFM and Wikipedia. As we show later in the results section, this inconsistency is due to the strong performance dependence of the Positive Only baselines on the underlying graph structure.
Figure 3. On unseen edges, the augmented algorithms (LINE II and node2vec II) yield transductive precisions comparable to that of the original algorithms, exceeding in certain cases. This improvement is seen most clearly in the cases of the Protein-Protein, Wikipedia, and LastFM graphs. The Positive Only Normalized baseline performs quite well but the performance relative to the baseline is not consistent, such as in the case of Protein-Protein and Wikipedia.
Figure 4. The above figure shows that the augmented algorithms, on average, perform less well on in-sample precision than the original LINE and node2vec algorithms. One explanation is that the dimension regularization prevents overfitting on the known edges. The Positive Only Normalized baseline performs well on in-sample precision because the algorithm exclusively trains on edge data and the normalization prevents the embeddings from collapsing.
Q2: How much of a runtime speedup does dimension regularization provide over SGNS?
Figure 5 shows that for both LINE and node2vec, replacing SGNS with dimension regularization significantly reduces runtime while preserving transductive precision. On average, replacing SGNS with dimension regularization decreased the runtime by $64$% for LINE and $83$% for node2vec. The light orange points in Figure 5 show the runtime and downstream performance for the augmented algorithms across all hyperparameter configurations while the dark orange points denote the best-performing configurations. As expected, completely removing the negative function results in the lowest runtime; thus the Positive Only baselines serve as lower bounds on the augmented algorithms’ runtimes. For the node2vec runtime comparison, we use code provided in Grover and Leskovec (2016) to generate the random walks and implement batched gradient descent in Python, as detailed in Algorithm 1 to create a level playing field with node2vec II. In practice, an optimized C implementation of our method would be even faster, much like node2vec’s speedup of the gradient descent step achieved via the gensim package. We include additional implementation details in Appendix B.2.
Figure 5. The above figure shows the training time and link-prediction performance for each algorithm on each graph. Compared to both vanilla LINE and node2vec (grey circle), our augmentation (orange star) maintains transductive precision while greatly reducing runtime. The results for our augmentation correspond with the optimal hyperparameter configuration; In Appendix B.2 we include runtime results for all configurations as well as details of our runtime evaluation methodology.
Q3: For which types of graphs does dimension regularization outperform all baselines?
We further investigate the limits of completely removing the negative function altogether and show that the inconsistency of the Positive Only baselines stems from its performance dependence on graph connectivity. We hypothesize that when the graph has low connectivity the initialization randomness allows nodes to be embedded near neighbors without collectively collapsing. At high connectivity however, collapse occurs.
We construct an experiment in which we measure transductive precision as the between-block edge probability in a Stochastic Block Model (SBM) is gradually increased. We define a two-block SBM where $n=200$ and the within-block edge probability $p$ is $0.2$. Then, we gradually increase the between-group edge probability $q$ until $p=q$.
Figure 6 shows that for low values of $q$ relative to $p$, all of the algorithms are able to to outperform a random baseline where top-$k$ predictions are uniform samples over $V$. When $q/p$ and graph connectivity is low, the blocks are more distinguishable and link-prediction is an easier task. However, as $q/p$ increases, the Positive Only performance quickly drops off whereas the augmented algorithm performance is more robust. We conclude that as graph connectivity increases, the Positive Only baseline is more prone to collapse. We do see that normalization does mitigate the collapse. Finally, as expected, when $p=q$, none of the embeddings outperform the random baseline as the SBM becomes an Erdös-Rényi graph.
Figure 6. While removing node repulsion altogether occasionally performed well for the real-world graphs, the Positive Only baselines rely on low graph connectivity to perform well. We construct a two-block SBM experiment in which the ratio of the between-block edge probability $q$ and the within-block probability $p$ is gradually increased. The figure above shows transductive precision relative to a uniform random baseline. When $q/p$ is low, the blocks are well-separated and link-prediction performance is high whereas when $p=q$, we expect no difference between any algorithm and the uniform-random baseline. In between, the augmented algorithm is most robust to increases in graph connectivity.
In Appendix C, we provide more fine-grained empirical evaluation which supplements the results shown here. Namely, analyze whether replacing SGNS with dimensionality reduction benefits or harms specific types of nodes and show that there are not discernable correlations between node feature and link-prediction performance changes. We also demonstrate that our regularization is capable of re-weighting nodes, just as in SGNS, as discussed in (15).
5. Related Works
In this section, we review the popular use of SGNS within graph embeddings and distinguish SGNS from the recent body of literature on negative sampling for self-supervised learning. Our algorithm augmentation is similar in spirit to the growing body of literature on non-contrastive learning, which we also review below.
5.1. Skip-Gram Negative Sampling
SGNS was introduced in word2vec by Mikolov et al. (2013) as an efficient method for learning word embeddings. While the softmax normalization constant is costly to optimize, Mikolov et al. (2013) modeled SGNS after Noise Contrastive Estimation (NCE) which learns to separate positive samples from samples drawn from a noise distribution. SGNS has since been adopted for graph representation learning where it is utilized in both unsupervised (Grover and Leskovec, 2016; Tang et al., 2015; Perozzi et al., 2014) and supervised skip-gram models (Yang et al., 2016).
In practice, the negative samples are drawn proportional to some function of node degrees, where many works raise the degree to the $3/4$ power (Grover and Leskovec, 2016). Yang et al. (Yang et al., 2020) study the proper noise distribution in the graph setting and propose sampling negative samples to balance downstream performance while minimizing the Noise Contrastive Estimation loss. This results in negative sample distributions that are positively but sub-linearly correlated with the positive sample distribution.
At the same time, there are many known limitations of SGNS. Rudolph et al. (2016) place SGNS embeddings within the framework of Exponential Family Embeddings and note that SGNS downweights the magnitude of the negative update and leads to biased embeddings, relative the gradients of the non-sampled objective. Second, Davison and Austern (2023) examine the limiting distribution of embeddings learned via SGNS and show that the distribution decouples from the true sampling distribution in the limit. Last, it has also been shown that the optimal noise distribution and the corresponding parameters can vary by dataset (Yang et al., 2020).
5.2. Non-contrastive Self-Supervised Learning
Energy Based Models in self-supervised learning are a unified framework for balancing similarity and dissimilarity (LeCun et al., 2007). As in our decomposition, energy-based models ensure that similar pairs have low energy and dissimilar pairs have high energy. Within energy-based models, there has been more focus across both vision and graph representation learning on contrastive models, which explicitly repel dissimilar pairs (In et al., 2023; Li et al., 2023; Yang et al., 2023; Zhang et al., 2023). However, given the computational complexities of pairwise contrast, there is a growing body of work on non-contrastive graph representation learning methods (Shiao et al., 2023b, a). Non-contrastive methods only use positive samples but then regularize to enforce dissimilarity. Shiao et al. (2023a) show that in certain cases, non-contrastive graph representation learning methods achieve comparable downstream performance while greatly reducing training time.
Recent works have shown that certain non-contrastive methods are simply efficient approaches to performing contrastive learning (Garrido et al., 2022). That said, these works specifically analyze the squared-loss term and there are no existing works, to our knowledge, establishing a connection between skip-gram loss and non-contrastive methods.

### v2

@@ 4.1. Instantiation for LINE and node2vec
LINE and node2vec differ in their implementation of the random-walk generation function. For LINE, the function simply returns the edge set $E$. For node2vec, the function returns the set of all node and neighbor pairs ($\left\{i,j|S_{ij}=1\right\}$) where neighbors of $i$ are nodes that co-occur in a biased-random-walk context window. The parameters $p,q$ control the bias of the random walk (Grover and Leskovec, 2016), interpolating between “outward” DFS-like walks and “inward” BFS-like walks. The default values for $p$ and $q$ are $1$; smaller values of $p$ increase the probability of backtracking to the previous node in the random walk and smaller values of $q$ encourage outward DFS-like walks. In Section 5.1, we discuss our hyperparameter optimization methodology for $p$ and $q$.
5. Experiments
To assess the efficacy of our dimension-regularization augmentation framework for capturing topological structure, we evaluate the link-prediction performance of the node2vec and LINE instantiations. Our results show that replacing SGNS with dimension-mean regularization reduces both training time and GPU memory usage while preserving link-prediction performance. Moreover, for real-world networks, we show that applying no repulsion at all, a special case of our framework, nearly always outperforms LINE and reduces training time by $71.7\%$ on average. We correlate the success of attraction-only models with local density, which is common in real-world networks. However, more generally, in networks with local sparsity, repulsion is needed, and dimension regularization provides an efficient alternative to SGNS.
5.1. Methodology
Below, we summarize the data and experimental methodology used for our empirical evaluation. All of the experiments were executed on a machine with a single NVIDIA V100 32 GB GPU. Our code is available at: https://doi.org/10.5281/zenodo.15557847. More details are provided in Appendix B.
Data.
We utilize the popular link-prediction benchmarks of Cora, CiteSeer, and PubMed (Yang et al., 2016) as well as four OGB datasets to demonstrate the scalability of our augmentation: ogbl-collab, ogbl-ppa, ogbl-citation2, ogbl-vessel  (Hu et al., 2020). We uniformly-randomly split the edge sets for Cora, CiteSeer, and PubMed into training/validation/test sets following standard 70/10/20 splits. For the OGB datasets, we use the provided edge splits. Table 2 summarizes the structural properties of the datasets, showing that the datasets span two types of densities: global density, measured by edge density, and local density, measured by the average clustering coefficient (“CC”) (Watts and Strogatz, 1998).
Table 2. Summary of datasets with structural properties. “CC” refers to the average clustering coefficient.
Dataset Nodes Edges CC Edge Density
Cora $2.7\times 10^{3}$ $1.2\times 10^{4}$ 0.24 $1.6\times 10^{-3}$
CiteSeer $3.3\times 10^{3}$ $1.0\times 10^{4}$ 0.14 $9.0\times 10^{-4}$
PubMed $2.0\times 10^{4}$ $9.8\times 10^{4}$ 0.06 $2.5\times 10^{-4}$
ogbl-collab $2.4\times 10^{5}$ $1.3\times 10^{6}$ 0.73 $2.4\times 10^{-5}$
ogbl-ppa $5.8\times 10^{5}$ $3.6\times 10^{7}$ 0.22 $1.1\times 10^{-4}$
ogbl-citation2 $2.9\times 10^{6}$ $3.1\times 10^{7}$ 0.09 $3.6\times 10^{-6}$
ogbl-vessel $3.5\times 10^{6}$ $5.9\times 10^{6}$ 0.01 $4.7\times 10^{-7}$
Models.
For each of node2vec and LINE, we instantiate three model variants: variant I is vanilla node2vec/LINE; variant II0 is a special instance of our framework in which no repulsion is applied at all ($n_{\text{negative}}$ is set to be larger than the number of batches); and variant II is our augmented model in which dimension regularization is applied at least once per epoch. The variant II0 is trained for only 2 epochs to prevent collapse. For completeness, we also instantiate variants of the vanilla algorithms and our augmentation in which node repulsion is proportional to $\text{degree}^{\alpha}$ as discussed in Section 3.2.1; these results are comparable to the results for I and II and included in Appendix C. In all cases, we set the embedding dimension $d=128$.
For link prediction, the embeddings for each model are used to train a multi-layer perceptron (MLP) edge classifier. A MLP is trained for each combination of dataset and model variant (I, II0, II). The input to MLP is the concatenation of two node embeddings and the output is an edge prediction score $\in[0,1]$. The time and memory values in the results correspond only to the resources used to train the embeddings and exclude the MLP training.
Hyperparameter optimization.
For each dataset and SG model, we optimize the following hyperparameters: the learning rate $\eta$, the node2vec random-walk parameters $p$ and $q$, as well as $n_{\text{negative}}$ and $\lambda$ from our augmentation. Details about the optimization process and the final values are included in Appendix B.
Metrics.
We report AUC-ROC, MRR, and Hits@$k$ evaluation metrics for link prediction. All metrics are evaluated at the node level and then averaged. To account for class imbalance when measuring AUC-ROC, we sample negative edges such that the number of positive and negative test edges is equal.
5.2. Results
5.2.1. Dimension regularization reduces training time and memory while preserving link-prediction performance.
Tables 3 and 4 show the training time, GPU memory usage, and AUC-ROC for vanilla and augmented node2vec and LINE. The training time is aggregated over multiple epochs, where the number of epochs is fixed for each graph and specified in Appendix B. The reported GPU memory is the maximum GPU memory allocated by PyTorch over the course of embedding training. The $\Delta(\%)$ columns under time and memory report the relative difference between II and I, which lower-bounds the efficiency improvement given that II0, a specific instance of our framework, is the most efficient variant.
For LINE, replacing SGNS with dimension regularization decreases training time by $11.4\%$, and link-prediction performance increases for all graphs, with AUC-ROC increasing by $0.05$ on average. Of note, LINE does not learn meaningful embeddings for the ogbl-vessel graph, likely due to the graph’s low edge density. For node2vec, training time reduces by $16.5\%$, on average. The AUC-ROC decreases for certain graphs, but, on average, the AUC-ROC changes by less than $0.05$. Meanwhile, GPU memory usage decreases by $30.7\%$ on average. In cases where GPU memory is constrained, our augmentation can enable node2vec training for graphs that were previously infeasible to train.
Our augmentation does not reduce memory usage for LINE because we obtain the best performance across all models when using small batch sizes; with small batches, even the vanilla algorithm has low GPU memory usage, and negative sampling contributes marginally to memory usage. We include the performance according to ranking metrics in Appendix C, which shows similar patterns as the AUC-ROC results.
Table 3. For LINE, our dimension-regularization augmentation framework decreases training time by $11.4\%$ on average. The differences in link prediction performance (as measured by AUC-ROC) are negligible (±$\pm 0.05$ on average). We use small batch sizes for LINE; thus, memory use is not affected. The column I refers to vanilla LINE. The column II refers to augmented LINE where dimension regularization is applied at least once per epoch. The column II0 refers to a special case of our framework where no repulsion is applied. For training time and memory, the $\Delta$ columns report the relative difference between II and I, which lower-bounds the efficiency gain of our framework as II0 is the most efficient.
Dataset Time (min) Memory (GB) AUC-ROC
I II0 II $\Delta(\%)$ I II0 II $\Delta(\%)$ I II0 II
CiteSeer 0.05 ¡0.01 0.04 -20.00% 0.02 0.02 0.02 0.00% 0.6 0.65 0.67
Cora 0.06 0.01 0.05 -16.67% 0.02 0.02 0.02 0.00% 0.59 0.67 0.68
PubMed 0.55 0.05 0.45 -18.18% 0.06 0.06 0.06 0.00% 0.7 0.72 0.75
ogbl-collab 0.77 0.25 0.65 -15.58% 0.64 0.64 0.64 0.00% 0.72 0.78 0.8
ogbl-ppa 16.23 5.12 15.43 -4.93% 2.07 2.07 2.07 0.00% 0.9 0.91 0.78
ogbl-citation2 48.51 25.19 47.4 -2.29% 8.02 8.02 8.02 0.00% 0.61 0.65 0.62
ogbl-vessel 9.39 4.9 9.2 -2.02% 9.19 9.19 9.19 0.00% 0.5 0.5 0.5
Table 4. For node2vec, our augmentation framework reduces training time by $16.5\%$ and memory by $30.7\%$ on average. The differences in link prediction performance (as measured by AUC-ROC) are negligible (±$\pm 0.05$ on average). The columns are defined analogously as in Table 3.
Dataset Time (min) Memory (GB) AUC-ROC
I II0 II $\Delta(\%)$ I II0 II $\Delta(\%)$ I II0 II
CiteSeer 0.53 0.03 0.42 -20.75% 3.12 2.09 2.09 -33.01% 0.76 0.78 0.77
Cora 0.44 0.03 0.36 -18.18% 3.12 2.08 2.08 -33.33% 0.87 0.66 0.8
PubMed 2.61 0.19 2.00 -23.37% 3.16 2.12 2.12 -32.91% 0.91 0.76 0.81
ogbl-collab 4.31 1.48 3.88 -9.98% 13.25 9.00 9.00 -32.08% 0.94 0.93 0.96
ogbl-ppa 24.72 9.23 23.43 -5.22% 27.24 18.74 18.74 -31.20% 1.00 0.81 0.97
ogbl-citation2 35.75 20.5 28.07 -21.48% 31.98 23.49 23.49 -26.55% 0.98 0.81 0.95
ogbl-vessel 51.69 27.93 42.99 -16.83% 32.84 24.35 24.35 -25.85% 0.96 0.83 0.87
5.2.2. Removing repulsion altogether reduces training time by over $70\%$ and improves LINE’s performance, especially for locally dense graphs.
Interestingly, we observe that the special case of our algorithm augmentation that removes repulsion altogether (II0) outperforms vanilla LINE for all seven graphs and significantly reduces training time. The top of Figure 3 shows that removing repulsion reduces training time by at least $40\%$. Meanwhile, the bottom of the figure correlates the gain in LINE’s performance with local edge density, as measured by the average clustering coefficient. The trend is most noticeable in the case of ogbl-collab, which has the highest clustering coefficient and reflects increases, relative to vanilla LINE, in all three metrics when repulsion is removed.
Intuitively, removing repulsion is effective in globally sparse but locally dense graphs, which characterizes many real-world networks (Newman, 2003), because the global sparsity hinders dimensional collapse while the local density ensures that node attraction brings nodes near hidden neighbors. While we would theoretically expect the embeddings to collapse even in sparse networks, in our experiments, we use the Adam optimizer, which decays the learning rate, and we train for only $2$ epochs in the case of II0. On the other hand, removing repulsion is less effective for node2vec because the random walks effectively increase global density–and thus, the likelihood of collapse–by creating positive pairs among not just connected nodes but also nodes within the same context window.
Figure 3. We find that in practice, removing repulsion altogether significantly reduces training time, as shown in the top figure, and can even improve performance. Specifically, removing repulsion performs especially well for real-world graphs that are globally sparse but locally dense (high clustering coefficient). In the bottom figure, we show the change in performance when repulsion is removed relative to vanilla LINE. For almost all graphs, AUC-ROC, MRR, and Hits@$k$ all increase; the increase is most prominent in the case of ogbl-collab, which has high local density.
@@ I II0 II $\Delta(\%)$ I II0 II $\Delta(\%)$ I II0 II
[Performance of removing repulsion completely]There are two plots. The top is a grouped bar chart with graph dataset names on the x-axis. For each dataset, there are two bars: one for the time decrease when removing repulsion from LINE and another for node2vec. The bottom plot contains three scatter plots, one for each of: AUC-ROC, MRR, and Hits@100. In each scatter plot, a point is a dataset, the x-axis is the average clustering coefficient, and the y-axis is the change in the metric value when repulsion is removed for LINE.
5.2.3. More generally, repulsion is needed, and dimension regularization provides a scalable solution.
To generalize beyond the seven real-world graphs, we examine the performance of our augmentation on two-block SBM graphs in which we toggle local density. Figure 4 shows that when the within-block edge probability is much greater than the between-block probability, all variants perform well. However, as local density decreases and the boundary between the two blocks erodes, repulsion is needed as indicated by the gap between I and II0; the figure shows that dimension regularization (II) is an effective repulsion mechanism. In LINE, dimension regularization outperforms all variants, and in node2vec, dimension regularization interpolates between the vanilla and repulsion-less variants. For LINE, the performance dip for II0 indicates that performance may be unstable when repulsion is removed, likely due to the possibility of embedding collapse.
[…]
[Removing repulsion for SBMs]The figure consists of two line charts. The first line chart shows the performance of each variant of LINE (AUC-ROC on the y-axis) as a function of the within-block / between-block edge-probability ratio. There is a line for each of the three variants of LINE. The right chart is the same, but for node2vec.
Figure 4. We show that as we decrease local density in a two-block SBM, the need for repulsion increases and dimension regularization provides an effective repulsion mechanism. In both figures, lower values on the x-axis imply fewer edges within blocks (lower local density) and more edges between blocks. When local density is high, the blocks are easily distinguishable and all variants perform well, but as local density decreases, repulsion is needed to separate the blocks and the performance of the repulsion-less special case (II0) erodes. Dimension regularization (II) performs the best in LINE and interpolates between the vanilla and repulsion-less variants in node2vec.
6. Discussion
In this section, we discuss the broader applicability of our theoretical result as well as the limitations of our dimension-regularization meta-algorithm.
Broader Applicability of Theoretical Result
While we focus on the skip-gram loss in this work, our theoretical contribution connecting skip-gram repulsion and dimension-mean regularization, as detailed in Section 3.2, is applicable beyond skip-gram. First, observe that the skip-gram loss we introduce in Eq. (8) is a generic cross-entropy logistic pair-classifier loss. Though we focus on the setting where positive and negative pairs are defined via the skip-gram model, dimension-mean regularization can be efficiently used for repulsion in Eq. (8) regardless of how positive and negative pairs are defined, provided the number of positive pairs is $o(n^{2})$, where $n$ is the vocabulary size. Second, dimension-mean regularization is also an efficiency bypass for certain formulations of Noise Contrastive Estimation (NCE) (Gutmann and Hyvärinen, 2012). Past works, such as the similarity-based graph embedding algorithm VERSE (Tsitsulin et al., 2018), have reformulated NCE in terms of negative sampling, using the exact SGNS loss in Eq. (9); for these negative-sampling-based formulations of NCE, the dimension-mean regularizer is an efficient bypass of negative sampling, again provided the number of positive samples is $o(n^{2})$. These two applications of the dimension-mean regularizer as an efficient form of repulsion for cross-entropy and NCE fall directly from our theoretical analysis in Section 3.2.
Meta-algorithm Limitations
While the results in Section 5 demonstrate that our augmentation framework reduces training time and memory usage while preserving link-prediction performance, we note two limitations. First, our dimension-regularization meta-algorithm introduces two hyperparameters $\lambda$ and $n_{\text{negative}}$, which require tuning. Second, while link-prediction performance is preserved overall, there are graphs, such as Cora, for which performance decreased relative to node2vec. That said, these graphs still benefit from significant decreases in training time and memory.